OpenStudy (anonymous):

A flying saucer with space aliens landed in my front yard! After the entire crew had filed out of their spaceship, I counted their heads and feet. I saw a total of 91 heads and 124 feet. it was clear that their leader was a purple creature with three heads and four feet. the rest of the space aliens each had two heads. the green ones had two feet and the blue ones had three feet. How many of each type landed on my lawn?

OpenStudy (anonymous):

@e.mccormick

OpenStudy (anonymous):

@Hero

OpenStudy (mathmath333):

the total were a 91 heads and 124 feet.

OpenStudy (mathmath333):

there was 1 leader (purple color) having with 3 heads and 4 feets

OpenStudy (mathmath333):

so total remaining had (91-3)=88 heads and (124-4)=120 feets

OpenStudy (mathmath333):

and the rest had 2 heads so total number(except leader) with green and blue color will be 88 heads divided by 2 heads =\(\Large 44~~ heads\) as each of them had two heads

OpenStudy (mathmath333):

so the total number of green and blue aliens will be \(\large g+b=44\)

OpenStudy (mathmath333):

now it is given that the green ones had \(2\) feet and the blue ones had \(3\) feet. so total number of remaining feets will be \(\large 2g+3b=120\)

OpenStudy (anonymous):

you need help

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} \normalsize \text{yes it will be}\hspace{.33em}\\~\\ g+b=44\hspace{.33em}\\~\\ \implies 2g+2b=88\hspace{.33em}\\~\\~\\ 2g+3b=120\hspace{.33em}\\~\\ \implies 2g+2b+b=120\hspace{.33em}\\~\\ \implies 88+b=120\hspace{.33em}\\~\\ b=32,g=12 \end{align}}\)

OpenStudy (mathmath333):

\(\huge \color{black}{\begin{align} \normalsize \text{so}\hspace{.33em}\\~\\ \color{purple}{\text{purple leader=1}}\hspace{.33em}\\~\\ \color{green}{\text{green aliens =12}}\hspace{.33em}\\~\\ \color{blue}{\text{blue aliens =32}}\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

@mathmath333 thank you